OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations auditioned today? 1300 1120 708 412

OpenStudy (anonymous):

I think the answer is 412

OpenStudy (anonymous):

x+y=1120 Eliminate the y: y=1120-x x=first step and y=second step The second equation is: 2x-2(1120-x)=1300 Solve: 5x-2240=1300 5x=3540 x=708 Now we would like to know the amountof the second step auditionerwhich is equal to y So plug in 708 into the equation of y: y=(1120-708) y=412

OpenStudy (anonymous):

If something isn't clear just ask me to explain